OpenStudy (anonymous):

How many photons are produced in a laser pulse of 0.548 J at 669 nm?

OpenStudy (anonymous):

Number of photons = 0.548 J / Energy of each photon Energy of each photon = Planck's Constant * frequency frequency = speed of light / wavelength Speed of light = 3.00 * 10^8 Wavelength = 669 nm or 669 * 10^-9 m Planck's Constant = 6.63*10^-34 J*s Calculate frequency, then energy of each photon, then number of photons.